D H-NH 53. In the Sherlock Holmes novel (1974) and movie The seven Per-cent solution a 7% solution of cocaine (C17H21NO4 303 g/mol) was used. Which of the following ratios would create a 7% solution? grams of cocaine in 100 mL of water B) 70 grams of cocaine in 1000 mL of water C) 35 grams of cocaine in 465 mL of water D) 21.21 grams of cocaine in 465 mL water E) none of the above 54, is a derivative of morphine that is as an or antitussive. It tastes very bitter. What is the pH of a 10mL solution containing 5 mg of pure codeine? 1.2 6 D)4 E) 9.6 55. A saturated solution of aspirin (HC9Hro4 180 contains 3gL. How would you change 1000mL of a saturated aspirin solution to a 0.01M solution? decreases to 900mL A) Add 250 mL of water B) heat the solution so the volume none of the above c) Add 0.6 g of aspirin D) add 667mL of water E)
Answer #1

53 )

answer : A ) 7 gm in 100 mL water

54 )

moles of codeine = 5 x 10^-3 / 300 = 1.67 x 10^-5

volume = 10 mL = 0.01 L

molarity = C = moles / volume = 1.67 x 10^-5 / 0.01

                = 1.67 x 10^-3 M

pKb of codeine = 6.05

pOH = 1/2 [pkb -logC]

pOH = 1/2 [6.05 - log 1.67 x 10^-3]

pOH = 4.41

pH + POH = 14

pH = 9.6

answer : E ) 9.6